How much energy does a photon with a wavelength equal to 350 nm have?

1. 2.32
Answer #1

The photon energy (E) is given by,

E = h*c/lambda = 5.67 E -19 J

where
h = 6.63 E-34 Js (Planck's constant)
c = 3.00 E+08 m/s (speed of light)
lambda = 350 E-09 m (wavelength)
1 nm = 1E-09 m
Answer #2

You can solve for energy with E = hv and c = wavelength * v

You need to subititute v with the second equation so that you can incorperate the wavelength into the energy equation.

So

V = c/wavelength

and thus: E = h(c/wavelength). Remember to convert nm (for the wavelength) to m as the speed of light is in meters.

5.68 * 10^-19 J

so answer is 4